sana purchases 3 lbs of mixed nuts and 89 cents per pound . hina purchases 2 lbs of peanuts at 49 cents per pound. what is the average price per pound for the mixture of both purchases ?
Answer by stanbon(75887)

sana purchases 3 lbs of mixed nuts and 89 cents per pound . hina purchases 2 lbs of peanuts at 49 cents per pound. what is the average price per pound for the mixture of both purchases ?
--------
Average = value/weight
-----
Value = 3*89 + 2*49 = 365 cents
Weight = 3 + 2 = 5 lbs
---
Average = 365/5 = 73 cents per pound
==========
Cheers,
Stan H.